OpenStudy (anonymous):

solve by substitution method 8x-3y=-94 2x+22=y

OpenStudy (anonymous):

8x-3y=-94 2x+22=y arrange them first 8x-3y=-94 2x-y=-22 multiply both sides by -4 then add ----------- 8x-3y=-94 -8x+4y=+88 ------------ 0 +y= -6 ans subs this to eq 2, 2x-y=-22 2x-(-6)=-22 2x+6=-22 2x=-22-6=-28 x=-28/2=-14 ans (-14,-6) ans